OpenStudy (anonymous):

A Web music store offers two versions of a popular song. The size of the standard version is 2.1 megabytes (MB). The size of the high-quality version is 4.9 MB. Yesterday, there were 770 downloads of the song, for a total download size of 2345 MB. How many downloads of the high-quality version were there?

OpenStudy (anonymous):

(2345-2.1s)/4.9=h

OpenStudy (anonymous):

You need to do a system of equations: 770=h+s and 2.1s+4.9h=2345

OpenStudy (anonymous):

standard=s, HQ=h s+h=770, 2.1s+4.9h=2345

OpenStudy (anonymous):

from there?

OpenStudy (anonymous):

i can substitute into the first equation?

OpenStudy (anonymous):

now solve for a variable, like i did above. "(2345-2.1s)/4.9=h" substitute this for h into the other equation so you have 770=h+s now: 770= [(2345-2.1s)/4.9]+s

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

to find s then substitute s into the second equation?

OpenStudy (anonymous):

once you have h you can substitute it into either. For this one the "770=h+s" would be a lot less messy

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

working...

OpenStudy (anonymous):

My suggestion is the elimination method. Multiply the first equation by -2.1 You have -2.1s - 2.1h=-1617. Add the two equations together. 2.8h=728 divide both sides by 2.8 h=260... Now just substitute 260 for h in the first equation and solve for s.

OpenStudy (anonymous):

just need the HQ # ^^)